### v5

@@ 5.1 Feasibility of Different Models
GPT-4-turbo can exploit up to 66% of the benchmark test-cases without human guidance. Taking into account high-level guidance, the success rates increase to 83%. This is comparable to the human pen-tester baseline that achieved 75% without hints or 92% with human-level guidance.
GPT-3.5-turbo fared worse being able only to achieve 16% success rates without guidance. Human-level guidance increased success rates to 50%. As gpt-3.5-turbo costs 20 times less than gpt-4-turbo, this might be an acceptable economic trade-off.
@@
Feasibility of Vulnerability Classes. Looking from the vulnerability class perspective, file-based exploits were well handled, information-disclosure based exploits needed human-level guidance for both LLMs and human ethical hackers, and multi-step cron attacks were hard for both LLMs and human operators.
5.2 Using State to Aggregate History
Results when using a LLM to summarize the current LLM’s world view into a compact state, and subsequently replacing history with that state, were surprising. When using less expressive LLMs, such as gpt-3.5-turbo or Llama3, success rates stagnated or even degraded. When using gpt-4-turbo for updating the state, success rates increased by 100% when unaided, and 25% when using human-level hints. Qualitative analysis indicates that this increase is due to gpt-4-turbo reflecting upon its existing knowledge of the target system and not only creating a new fact list (worldview) but also including potential attack vectors for subsequent rounds. This indicates the benefits of using the Reflection pattern.
@@ 5.2 Using State to Aggregate History
During the evaluation, one drawback was identified: the update-state prompts took significantly longer than the next-cmd prompts even when the latter included the history. Using GPT-4, the update-state queries took $13.4$ times longer than the next-cmd queries ($19.89s$ vs. $1.48s$ on average). Another problem is OpenAI’s asymmetric pricing of tokens: output tokens, e.g., tokens needed for updating the state cost thrice as much as input tokens, thus making state-processing potentially cost-ineffective.
5.3 Impact of Guidance
(a) GPT-3.5-turbo with maximum context size 8k.
(b) GPT-4-turbo with maximum context size 8k.
Figure 6: Context Token Usage by different models. Colors indicate different test-cases and are identical in both graphs.
We implemented two guidance mechanisms: on one hand, a single hint per test case akin to hints given by high-level LLM Planner modules, was implemented. On the other hand, influenced by human pen-tester behavior observed during our human baseline evaluation, we implemented autonomous guidance based upon LLM-analysis of initial enumeration tool results.
High-level guidance consistently improved success rates. When using smaller models, e.g., gpt-3.5-turbo or Llama3, they were mandatory to achieve acceptable access rates, e.g., gpt-3.5-turbo’s success rate increased from 15% to 50%. When using gpt-4-turbo, guidance increased success rates from 33% to 66%.
Enumeration-based automatic guidance had less impact. They slightly improved success rates, typically allowing LLMs to achieve one additional successful test-case. Qualitative analysis showed that human hints help human hackers and automations to think “outside the box” and pursue new attack vectors, e.g., search for passwords in files, while enumeration-based hints kept the pen-testers and automations “within the box”.
5.4 Impact of Context-Size
The maximum available context size highly depends on the respective model. In addition to the maximum token count, different model families use different tokenizers thus making context sizes not directly comparable between LLM families. For example, the amount of tokens generated by OpenAI’s tokenizer (used by GPT-3.5-turbo and GPT-4-turbo) was smaller than the amount produced by the llama one. Another difference is what data is counted towards the context size limit. For example, OpenAI-based models only count input tokens, i.e., the used prompt, while while Llama3-based models count both input and output data, i.e., the used prompt plus the generated answer.
@@ 5.4 Impact of Context-Size
We used a context size of 8192 as an initial limit. This context size is supported by GPT-3.5-turbo, GPT-4-turbo as well as by the different Llama3 models. In addition, using a smaller context size should reduce computation time and directly impact occurring query costs.
@@
Figure 6 shows the context usage counts during different runs utilizing OpenAI models, indicating that GPT-3.5-turbo is using up context quicker than gpt-4-turbo [^31] Please note, that we are adding both prompt and answer token counts, so the sum can be larger than the input limit of $8192$ tokens.. When looking at the executed commands, GPT-3.5-turbo is filling up the context size with output of “broad” commands such as “ps aux” or “find / -type f” commands while GPT-4 executes more targeted commands whose results require little context size.
@@
Figure 7: Context Token Usage by different models. Colors indicate different test-cases and are identical in both graphs.
@@
GPT-4-turbo’s large context size of 128k allowed us to utilize in-context learning. We used roughly 50% of the available context size to include background hacking information extracted from hacktricks. Adding this hacking background did not yield better results, indicating no benefit over the information inherently stored within the LLMs themselves. We assume that this background information would help smaller models as they store less information within their model weights. Alas, Llama3’s small context size of 8k prevented empirically testing this assumption.
@@
Using in-context learning substantially increases benchmarking costs. Gpt-4-turbo is currently billed \$10 for one million input tokens, thus including background information adds costs of \$$\$0.67$ per gpt-4-turbo prompt called. In our worst-case scenario of a benchmark run using 12 test cases with 60 rounds per test case, including hacking background would add costs of \$$\$482.4$ per benchmark run in addition to the actual prompt costs [^33] Please note that recent changes in OpenAI’s and Anthrophic’s could based offering implement prefix prompt caching, reducing the cost of reused initial prompts by $50\%$ and $90\%$ respectively..
@@ 6 Discussion
While metrics provide a quantitative overview of the efficacy of wintermute, we also inspected the quality of the generated Linux privilege-escalation commands based on data collected during benchmarking. We discuss notions of causality and common-sense in multi-step exploits and provide a comparison to commands and strategies typically seen by human pen-testers in similar situations.
@@ 6.1 LLM-Generated Commands
Successful privilege escalation depends upon the quality of LLM-provided commands. OpenAI-based models can consistently provide valid Linux commands while Llama3-based models struggled. As shown in Figure 8, generated Llama3-70b commands often contained syntax errors or Llama3 interleaved comments into commands, thus making them invalid. Llama3-8b was often not able to correctly call capabilities but hallucinated new capabilities such as exec_cmd, exec_find, or exec_cat [^34] A workaround was added to wintermuteto detect these hallucinations and execute the intended commands..
@@
Figure 8: Examples of overly complex or faulty commands generated by LLMs.
@@ 6.2 Reacting to System Responses
Thinking inside the box. Although LLMs were able to identify potential passwords in configuration files, e.g., for database users, they did not test those for password-reuse, i.e., if the root user was reusing one service account password. This “out-of-the-box thinking” occurs commonly during pen-testing. We assume that nothing in the model was able to statistically map those occurrences to a privilege escalation path while humans were commonly able to do this.
Ignoring Responses. All tested LLMs were repeating almost identical commands and thus wasted rounds as well as resources. Occurrences included repeated enumeration commands (“sudo -l”, “cat /etc/passwd”), retesting the same credentials, or calling “find” for locating files. The latter was often called with syntactical variations while keeping the semantics of the operation the same, e.g., different order of parameters or using “-perm u=s” instead of “-perm /4000”. This indicates that LLMs were acting as stochastic parrots without deeper understanding of the uttered commands’ semantics.
Related to both this and the next topic, LLMs often threw away potential error messages by appending “2>/dev/null” to generated commands.
@@
Another example of this is LLMs ignoring direct error messages, e.g., GPT-3.5-turbo tried to keep using sudo even when each invocation returned an error that the user is not included in the sudoers file and thus now allowed to use sudo.

### v6

@@ 5.1 Feasibility of Different Models
RQ1: What is the efficacy of LLMs in performing autonomous Linux privilege escalation attacks? Answer: We show that small language models such as Llama3-8b are currently not capable for autonomous penetration-testing. Larger models, such as GPT-3.5-Turbo and Llama3-70b are capable of matching the efficacy of traditional automated privilege-escalation tooling (success rates ranging from 8 to 40%). GPT-4-Turbo is able to surpass traditional tooling and its success rates (33-66%) are similar to success rates of human professional penetration-testers (75%) while their imposed costs are competitive to the cost of professional penetration-tests (Section 5.4).
GPT-4-Turbo can exploit up to 66% of the benchmark test-cases without providing high-level test-case specific hints. Taking into account high-level guidance, the success rates increase to 83%. This is comparable to the human penetration-tester baseline that achieved 75% without hints or 92% with high-level guidance.
GPT-3.5-Turbo fared worse being able only to achieve 16% success rates without providing high-level guidance. high-level guidance increased success rates to 50%. As GPT-3.5-Turbo costs 20 times less than GPT-4-Turbo, this might be an acceptable economic trade-off.
@@
Feasibility of Vulnerability Classes. Looking from the vulnerability class perspective, file-based exploits were well handled, information-disclosure based exploits needed high-level guidance for both LLMs and human ethical hackers, and multi-step cron attacks were hard for both LLMs and human operators.
5.1.1 Using State to Aggregate History
Results when using a LLM to summarize the current LLM’s world view into a compact state, and subsequently replacing history with that state, were surprising. When using less expressive LLMs, such as GPT-3.5-Turbo or Llama3, success rates stagnated or even degraded. When using GPT-4-Turbo for updating the state, success rates increased by 100% when unaided, and 25% when using high-level hints. Qualitative analysis indicates that this increase is due to GPT-4-Turbo reflecting upon its existing knowledge of the target system and not only creating a new fact list (worldview) but also including potential attack vectors for subsequent rounds. This indicates the benefits of using the Reflection pattern (Renze and Guven, 2024).
@@ 5.1.1 Using State to Aggregate History
During the evaluation, one drawback was identified: the update-state prompts took significantly longer than the next-cmd prompts even when the latter included the history. Using GPT-4-Turbo, the update-state queries took $13.4$ times longer than the next-cmd queries ($19.89s$ vs. $1.48s$ on average). Another problem is OpenAI’s asymmetric pricing of tokens: output tokens, e.g., tokens needed for updating the state cost thrice as much as input tokens, thus making state-processing potentially cost-ineffective.
5.2 Impact of Context Management Strategies and Context-Size
RQ2: How do various context management strategies and context sizes impact the efficacy and efficiency of LLM-driven privilege escalation agents? Answer: Our results show that using GPT-4-Turbo to compress history into a compact state can increase success rates by 100% unaided and 25% with high-level hints, indicating the benefits of the Reflection pattern. We also explore the efficacy of adding background hacking information and the impact of larger execution histories, noting that background hacking information had a smaller impact with state-of-the-art models like GPT-4-Turbo, while increasing stored history significantly improved success rates. However, using larger context sizes and in-context learning substantially increased benchmarking costs.
The maximum available context size highly depends on the respective model. In addition to the maximum token count, different model families use different tokenizers thus making context sizes not directly comparable between LLM families. For example, the amount of tokens generated by OpenAI’s tokenizer (used by GPT-3.5-Turbo and GPT-4-Turbo) was smaller than the amount produced by the Llama3 tokenizer. Another difference is what data is counted towards the context size limit. For example, OpenAI-based models only count input tokens, i.e., the used prompt, while while Llama3-based models count both input and output data, i.e., the used prompt plus the generated answer.
@@ 5.2 Impact of Context Management Strategies and Context-Size
We used a context size of 8192 as an initial limit. This context size is supported by GPT-3.5-Turbo, GPT-4-Turbo as well as by the different Llama3 models. In addition, using a smaller context size should reduce computation time and directly impact occurring query costs. We reduced the input token size of Llama3-based models to 6k to allow for output tokens.
@@
Figure 6 shows the context usage counts during different runs utilizing OpenAI models, indicating that GPT-3.5-Turbo is using up context quicker than GPT-4-Turbo [^25] Please note, that we are adding both prompt and answer token counts, so the sum can be larger than the input limit of $8192$ tokens.. When looking at the executed commands, GPT-3.5-Turbo is filling up the context size with output of “broad” commands such as ps aux or find / -type f commands while GPT-4-Turbo executes more targeted commands whose results require little context size.
@@
Figure 5: Graph of accumulated context token usage over time (indicated by the rounds) for different LLMs. Colors indicate different test-cases and are identical in both graphs (also see Figure 6).
@@ 5.2.2 In-Context Learning (ICL)
GPT-4-Turbo’s large context size of 128k allowed us to utilize In-Context Learning (ICL). Based on prior research indicating that ICL typically rivals Retrieval Augmented Generation (RAG) performance for retrieval tasks (Lewis et al., 2020a; Lee et al., 2024; Li et al., 2024), we used ICL to test the efficacy of integrating external knowledge as a proxy for RAG-based approaches. We used roughly 50% of the available context size to include background hacking information extracted from hacktricks. Adding this hacking background did not yield better results, indicating no benefit over the information inherently stored within the LLMs themselves. We assume that this background information would help smaller models as they store less information within their model weights. Alas, Llama3’s small context size of 8k prevented empirically testing this assumption.
@@
Using in-context learning substantially increases benchmarking costs. GPT-4-Turbo is currently billed \$10 for one million input tokens, thus including background information adds costs of $\mathdollar 0.67$ per GPT-4-Turbo prompt called. In our worst-case scenario of a benchmark run using 12 test cases with 60 rounds per test case, including hacking background would add costs of $\mathdollar 482.4$ per benchmark run in addition to the actual prompt costs [^27] Please note that recent changes in OpenAI’s and Anthrophic’s could based offering implement prefix prompt caching, reducing the cost of reused initial prompts by $50\%$ and $90\%$ respectively..
One concern when using ICL is the LLM’s limited context size. Within our prototype, context is used by both the background information (used by ICL) and the accumulated command execution history which is truncated to fit the available context size. If ICL occupies too much context, storage for command history is reduced leading to the prototype discarding older parts of the execution history, potentially loosing information needed for successful exploitation.
Analysis of the included background knowledge has shown that its size is roughly 63k tokens (Section 3.5.3). When investigating the impact of ICL, we are using GPT-4-Turbo with its 128k context size, allowing for 65k tokens for history storage. Figure 5(a) visualizes the token usage of a standard GPT-4-Turbo test-run, showing that a single run used more than 20k tokens. This indicates that only during a test-case history truncation would occur and otherwise both history and background knowledge fit into the model’s context size.
5.3 Impact of Guidance
RQ3: To what extent do different guidance mechanisms influence the success rates of attack vectors by LLM-based privilege-escalation agents? Answer: We show that high-level guidance, similar to that provided by human penetration-testers or through privilege-escalation check-lists, can significantly increase success rates, for example, from 66% to 83% for GPT-4-Turbo, and from 16% to 50% for GPT-3.5-Turbo. The study also examined enumeration-based automated guidance, which improved success rates but was less effective than high-level guidance, partially because enumeration tools tend to “stay in the box”.
(a) GPT-3.5-Turbo with maximum context size 8k.
(b) GPT-4-Turbo with maximum context size 8k.
Figure 6: Graphs of accumulated context token usage over time (indicated by the rounds) for different LLMs. Both graphs use identical colors for identifying test-cases.
We implemented two guidance mechanisms (Section 3.5.2): on one hand, a single high-level hint was provided; on the other hand, we implemented autonomous guidance based on LLM-analysis of an initially executed traditional enumeration tool. The latter was influenced by observing our human baseline penetration-tester that typically executed an enumeration-script (lse.sh) before basing their next steps on the enumeration script’s output. Results of enumeration-based guidance can be seen in Table 8: the column Guidance is used to indicate if either a high-level hint(“h”) or enumeration-based guidance (“e”) was supplied to the prototype.
High-level guidance consistently improved success rates. When using smaller models, e.g., GPT-3.5-Turbo or Llama3, they were mandatory to achieve acceptable access rates, e.g., GPT-3.5-Turbo’s success rate increased from 16% to 50%. When using GPT-4-Turbo, guidance increased success rates from 33% to 66%. Enumeration-based automatic guidance had less impact. They slightly improved success rates, typically allowing LLMs to achieve one additional successful test-case, e.g., the success rates of GPT-3.5-Turbo increased from 16% to 24% while the success rates of GPT-4-Turbo increased from 33% to 40%.
We also investigated the use of multiple LLMs within a single prototype. For this, we evaluated a setup utilizing enumeration-based guidance in which GPT-4-Turbo was used to extract three hints from the output of the enumeration tool, and used GPT-3.5-Turbo to attempt privilege-escalation using these extracted hints. This uses the advanced but expensive GPT-4-Turbo for the initial analysis step, the faster and cheaper GPT-3.5-Turbo model is used for the more frequent execution steps. Using this combination increased the success rates from between 16% (no guidance) and 24% (enumeration analysis through GPT-3.5-Turbo) to 40%, matching the performance of using GPT-4-Turbo for both enumeration analysis and tool execution.
Qualitative analysis showed that high-level hints help human hackers and automations to think “outside the box” and pursue new attack vectors, e.g., search for passwords in files, while enumeration-based hints kept the pen-testers and automations “within the box”.
5.4 Cost Analysis
Table 9: Current costs of using different LLMs hosted at Provider. All costs are given in US\$ per million Input/Output tokens.
LLM Provider Cost/mToken Input Cost/mToken Output
Llama3-8b DeepInfra $\mathdollar 0.03$ $\mathdollar 0.06$
Llama3-70b DeepInfra $\mathdollar 0.30$ $\mathdollar 0.40$
GPT-3.5-Turbo OpenAI $\mathdollar 0.50$ $\mathdollar 1.50$
GPT-4-Turbo OpenAI $\mathdollar 10.00$ $\mathdollar 30.00$
We analyze the monetary impact of LLM-driven penetration-testing by calculating the costs (in US\$) that would have occurred if all tests had been performed during July, 2025. We use the amount of consumed tokens and use current pricing to calculate costs. For OpenAI-based models, we use their current token pricing; for locally-ran Llama3 models we use the pricing of a cloud-hosted Llama3 model on DeepInfra282828deepinfra.com to create comparable costs. The resulting costs are listed in Table 9.
We base the costs of our human baseline on their invested time (1 hour). indeed.com, a meta-search engine that aggregates job-postings, estimates the average salary of a penetration-tester at US\$ 53/hour. This fits with penetration-testing companies typically charging between US\$ 100–300/h for penetration tests. Based on this data, we chose a conservative penetration-testing rate of US\$ 159/hour (three times the salary of a penetration-tester).
Calculated cost-estimates for both LLM-driven prototypes and human penetration-testers are shown in Table 10. We differentiate between three use-cases for our analysis: first, we analyze the results from the point-of-view of a company that would hire professional penetration-testers for a single project and currently has no access to human professional penetration-testers or cannot afford their cost (baseline: Human Baseline: Cost of Pentest). Second, we analyze the results from the point-of-view of a penetration-testing company that wants to extend or replace its workforce with LLM-driven penetration-testers (baseline: Salary-based, as the cost of LLM-driven penetration-testing should not be larger than that of human penetration-testers). And finally, we analyze the potential of augmenting human penetration-testers with LLM-driven agents by analyzing the costs when using the emulated guidance: guided provided by our testbed.
The amount of solved test-cases, e.g., successfully exploited vulnerabilities, indicates that neither Llama3-based systems nor GPT-3.5-Turbo  are currently suitable to replace or augment human penetration-testers. This matches results indicates by prior research (Section 2.3). Given their low operating costs of \$0.005 to \$0.43 per successfully exploited vulnerability, we recommend further research into increasing their efficacy.
Next, we analyze the results of fully autonomous (not using high-level hints as guidance) prototype runs that achieve similar success levels as human penetration-testers. This effectively limits our model-selection to GPT-4-Turbo. Two prototype runs look promising: using GPT-4-Turbo with a context-size of 8k and the improved state-management mechanism (cost: \$1.54 per exploited vulnerability), and using GPT-4-Turbo with a context-size of 128k (cost: \$11.43 per exploited vulnerability). The former is suitable for both companies in need of a pentest (use-case 1) as well as for companies that want to replace existing penetration-testers (use-case 2). The latter would only be suitable for companies in need of pen-test without access to human penetration-testers as its costs is higher than the salary of a penetration-tester thus making it inefficient for penetration-testing companies to use this to replace existing penetration-testers.
For analysis of augmenting human penetration-testers we look at GPT-4-Turbo  using high-level guidance where the high-level hint is a standin for human-given tasks. When using GPT-4-Turbo  (8k) with state-management, high-level guidance (i.e., giving the prototype an area to investigate) roughly reduces the costs by 50%. While incorporating background hacking knowledge through In-Context Learning (ICL) allowed for high successful exploitation rates, its costs would surpass the cost of human penetration-testers and thus is inefficient to use.
Table 10: Cost Analysis of Human Baseline and LLM-driven Prototype Runs. For calculating costs, we are using token costs of hosted LLMs (July, 2025) and average penetration-tester salaries and penetration-test costs for our human baseline.
Query Response Cost (US\$)
Context
Memory
Guidance
solved
kTokens
Cost (US\$)
kTokens
Cost (US\$)
Total
per Vuln
Human Baseline
Salary-based 9 53 5.89
Cost of Pentest 9 159 17.67
Llama3-8b
8k hist. - 0 133.0 0.004 23.9 0.001 0.01 -
8k hist. hint 2 301.1 0.009 30.2 0.002 0.01 0.005
Llama3-70b-q4
8̃k hist. - 3 2778.5 0.834 15.3 0.006 0.84 0.28
8̃k state - 1 698.1 0.209 131.1 0.052 0.26 0.26
8̃k hist. hint 4 2106.5 0.632 24.1 0.01 0.64 0.16
GPT-3.5-Turbo
8k hist. - 2 3250.3 1.625 8.1 0.012 1.64 0.82
8k state - 2 630.9 0.315 60.6 0.091 0.41 0.205
8k hist enum 3 2540.0 1.27 8.8 0.013 1.28 0.43
GPT-4-Turbo (Pen-Tester Replacement)
4k hist. - 4 1133.8 11.338 6.8 0.205 11.54 2.89
4k state - 6 737.1 7.371 189.7 5.69 13.06 2.18
8k hist. - 4 2162.2 21.622 6.3 0.19 21.81 5.45
8k state - 8 724.9 7.249 169.1 5.073 12.32 1.54
8k hist. enum 5 1405.9 14.059 6.5 0.194 14.25 2.85
128k hist. - 8 9118.3 91.183 8.6 0.259 91.44 11.43
GPT-4-Turbo with hints (Pen-Tester Support)
8k state hint 10 451.2 4.512 111.5 3.346 7.86 0.79
128k (ht) hist. hint 9 25605.9 256.059 5.6 0.169 256.23 28.37
The table is grouped by the used LLM model family, e.g., GPT-4-Turbo or Llama3. For each executed prototype-run, we list the allowed maximum context size (Context), the used Memory technique (hist. identifies runs that used truncated memory, state denotes runs that used the summarized state technique), Guidance shows if hints were supplied to the LLM (either high-level hints or automated enumeration based hints), solved is the number of successfully exploited virtual machines, Query and Response highlights the token-cost for running the whole benchmark-set of 12 test-cases, both in kilo-Tokens as wel as in US\$. Finally, Cost lists the overall costs of a benchmark run in both US\$ as well as the average cost in US\$ per exploited virtual machine.
@@ 6 Discussion
While metrics provide a quantitative overview of the efficacy of hackingBuddyGPT, we also inspected the quality of the generated Linux privilege-escalation commands based on data collected during benchmarking. We discuss notions of causality and common-sense in multi-step exploits and provide a comparison to commands and strategies typically seen by human pen-testers in similar situations.
@@ 6.1 LLM-Generated Commands
Successful privilege escalation depends on the quality of LLM-provided commands. OpenAI-based models can consistently provide valid Linux commands while Llama3-based models struggled. As shown in Figure 7, generated Llama3-70b commands often contained syntax errors or Llama3 interleaved comments into commands, thus making them invalid. Llama3-8b was often not able to correctly call capabilities but hallucinated new capabilities such as exec_cmd, exec_find, or exec_cat. A workaround was added to hackingBuddyGPT to detect these hallucinations and execute the intended commands.
@@
Figure 7: Examples of overly complex or faulty commands generated by LLMs. Problems include incorrectly generated commands (wrong parameter order, including comments such as “,etc.”) as well as overly convoluted commands.
@@ 6.2 Reacting to System Responses
Thinking inside the box. Although LLMs were able to identify potential passwords in configuration files, e.g., for database users, they did not test those for password-reuse, i.e., if the root user was reusing one service account password. This “out-of-the-box thinking” occurs commonly during penetration-testing. We assume that nothing in the model was able to statistically map those occurrences to a privilege escalation path while humans were commonly able to do this.
Ignoring Responses. All tested LLMs were repeating almost identical commands and thus wasted rounds as well as resources. Occurrences included repeated enumeration commands (sudo -l, cat /etc/passwd), retesting the same credentials, or calling “find” for locating files. The latter was often called with syntactical variations while keeping the semantics of the operation the same, e.g., different order of parameters or using -perm u=s instead of -perm /4000. This indicates that LLMs were acting as stochastic parrots without deeper understanding of the uttered commands’ semantics.
Related to both this and the next topic, LLMs often threw away potential error messages by appending 2>/dev/null to generated commands.
@@
Another example of this is LLMs ignoring direct error messages, e.g., GPT-3.5-Turbo tried to keep using sudo even when each invocation returned an error that the user is not included in the /etc/sudoers file and thus now allowed to use sudo.